In a walk a thon to rais e money for charity, Elisa walked a certain distacne at 5mph and then jogged twice the distacne at 8mpn. Her total time walked and jogged was 2 hours and 15 minute. How far did Elisa get?
Answer by checkley71(8403)

D/5+2D/8=2.25 OR 18D/40=2.25 OR 18D=90 OR D=5 MILES.
THEREFORE 5/5=1 HOUR WALKING 2D/8 OR 10/8=1.25 HOURS OR 5 MILES +10 MILES =15 TOTAL MILES.